OpenStudy (anonymous):

3x^2=6 solutions? x-intercepts?

OpenStudy (anonymous):

\[-\sqrt{2},\sqrt{2}\] x-intercepts?

jhonyy9 (jhonyy9):

3x^2 =6 x^2 =2 x_1,2=+/- sqrt2

OpenStudy (anonymous):

x-intercept is an ordered pair